Day shift summary 26/dec/99

ERWIN@wishpa.physics.wisc.edu
Sun, 26 Dec 99 15:40 CDT

Shift Summary, Day shift, 26/Dec/99

On Shift: Vivian, Albert

Hours of beam available at all intensities: 7.15
Hours of beam available at nominal intensity: 7.15
Hours of E799 data (special runs not included): 6.10
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Pipeline errors. Beam down.

========================================================================
Runs ended this shift: 15292, 15293, 15294
Run 15295 BEAM_799 in progress

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: None

========================================================================
Overview of shift: For about an hour the SDAQ had red buttons on SDAQ,
PROCESSES, and TRD-GAS. A fresh start for SDAQ would not cure the
problem. Theo was called for help. "About once a month the 'Serious
Alarms' monitor needs to be rebooted. This can be done by choosing
an option on the monitor tool bar. The TRD-GAS system complains that
there is a 'broken pipe', so it must be rebooted too from the monitor
in rack 200 in far SW corner of RR area. SDAQ error message will say
TRD-GAS needs to be rebooted, but doesn't tell you that it can't be
done if the serious-alarms monitor needs to be rebooted.

Stubborn pipeline errors resisted reset, KFREND, and power cycling.
Bob T approved a fresh start, but Viv just started a new run. That
cured everything. May have lost about an hour of run time here.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15292 07:00 Peaceful run inherited at 7E12 from previous shift
15292 09:15 SDAQ goes bad with red buttons. Eventually Theo
nails "serious-alarms" monitor as needing a reboot.
After that TRD-GAS monitor could be successfully
rebooted.
15292 09:45 "Tapes full" complaint when colored bar graphs
indicate only about 60% full. Viv notes a couple of
individual tapes are nearly full. Without level 2
triggers running, tapes are not so uniformly filled.
Drive 3 on ktev1 seems to be empty.
15293 09:45 Load new tapes and start new run 15293
10:55 Incurable pipe line errors.
15294 11:15 Simply starting a new run fixed the problem.
(no fresh start)
11:30 MI PS "complications", investigating, no beam
11:54 Beam back
14:15 MI PS trips, no beam
14:40 MCR consulting with experts on PS trip. No estimate
14:40+ Vivian discovers that the event size being written
to KTeV3, Drive 3 is ZERO. This may explain partially
why that tape does not seem to fill up.
14:56 After consulting with Ed B., Viv stops run
and prepares drt_fresh_start. Also started
scanning the tape from last night to see if the
monitoring is really correct.
15:30 Still no beam. Fresh_start seems to have
fixed tape drive monitoring problem (?). To
be followed up by the able bodies on the next
shift.

========================================================================
Follow up Notes on Previous Shift
Pray for the recovery of the fruit.

========================================================================
Items to Follow up on Next Shift
Continue to worry about empty tape on ktev1, drive 3.